From the 1/14th Daily Journal Summaries for the month of April, 1968.

April 8-14, 1968

8 April 1968


GOLDEN DRAGONS deployed from Phu Cat AB to Polie Kleng Via (6) C-130 sorties. The 1/14 Infantry closed South of the Polei Kleng Airstrip and prepared for future operations.

Plans: Golden Dragons prepare for operations in MacArthur AO under present control of 3d Bde, 4th Infantry Division.

9 April 1968


GOLDEN DRAGONS conducted resupply at refit activities at Polie Kleng in preparation of future operations in the MacArthur AO. Resupply was conducted from Kontum to Polie Kleng by CH-47.

Plans: 1-14th Infantry will conduct Combat Assault to the AO North of Polie Kleng and conduct search and destroy operations secure the firebase and prepare it for future occupation. Bravo Company will Combat Assault to exploit an arclight Strike East fo the firebase.

10 April 1968


GOLDEN DRAGONS conducted multiple CA's after Airstrikes were placed on LZ No Slack. Charlie, Alpha, Reconnaissance platoon Combat Assaulted to the firebase sight. The LZ was Cold and elements of the TOC CP accompanied the Combat Assault with the Engineer Platoon.

Plans: GOLDEN DRAGONS continued construction on LZ No Slack and continue offensive operations and exploit the Arclight.

11 April 1968


GOLDEN DRAGONS Continued operations in their AO North of Polie Kleng and Southwest of Dak To on the Cu Tosung Mountains. Bravo Co. conducted a Combat Assault into the Southern portion of the Arclight and moved to the Northwest to exploit the strike. Battalion HVY mortar Plat was lifted to LZ No Slack (Blue). Alpha Company swept south from LZ No Slack (Red) and returned to night loger South of LZ No Slack with neg contact.

Plan: Delta Company will conduct a secure move to LZ No Slack (Red) and push to the Southeast to sweep the high ground East of LZ No Slack. CP(-) will be lifted into the firebase. Construction continues on LZ No Slack for employment of the artillery pieces.

12 April 1968


GOLDEN DRAGONS the morning of 12 April saw Co B moving west from LZ No Slack and Co D conducting a move from Ploei Kleng LZ to LZ No Slack. The Bn Hq finished movement to LZ No Slack from LZ Polei Kleng in the early afternoon. GEN Hickman arrived at LZ No Slack and was briefed by the S-3 at 1225. The situation remained quiet for all elements throughout the day and night.

13 April 1968


GOLDEN DRAGONS in the early morning hours, the 1st plat of Co A, the security unit around TOC, detected movement on the south side of the perimeter. No further movement was detected during the early morning hours. Co B ande D moved out further into the AO. Co C moved south towards LZ No Slack to take up a security mission on the south portion of the LZ. In the late afternoon Co D spotted 3 or 4 enemy and called artillery, with negative results. In the early evening C A detected movement to the Northeast, and are continuing to observe.

14 April 1968


GOLDEN DRAGONS No further reports were received from Co A on Hill 1018. Hq 3d Bde sent OPLAN 7-68 to the Bn during the mid-morning hours. BG Hickman arrived at LZ No Slack and was briefed by LTC BAll at 1230. At 1305 Co D detected and took under fire 2 enemy moving west to east, with negative results. Night found B and C Co's in the AO with A and C still securing LZ No Slack.

APRIL 15-21